# Data-Driven Distributed Optimization via Aggregative Tracking and Deep-Learning

### v2

@@ IV-B Boundary-Layer System Analysis
In this section, we analyze the so-called boundary-layer system associated to (18), i.e., the dynamics obtained by considering an arbitrarily fixed $(\tilde{\theta},\tilde{\mathrm{x}})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ into (18c)-(18d) written in the error coordinates
@@ IV-C Reduced system analysis
By using the accuracy characterization (6c) of $\theta^{\star}$ and since $x^{\star}$ is the solution to (1), we point out that the origin is an equilibrium point of (25). We start by analyzing the stability of the origin for the corresponding averaged system, i.e., the auxiliary scheme obtained by averaging the time-varying vector field driving (25) over an infinite time-horizon (see Appendix A-A). To this end, let $\tilde{\ell}_{\mathcal{K}}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ be
$\tilde{\ell}_{\mathcal{K}}(\tilde{x},\tilde{\theta})\coloneq\frac{1}{\mathcal{K}}\sum_{k=1}^{\mathcal{K}}\ell(\tilde{x}\!+\!x^{\star}\!+\!\mathrm{d}^{k}_{x},\mathbf{1}\sigma(\tilde{x}\!+\!x^{\star})\!+\!\mathrm{d}^{k}_{\sigma},\tilde{\theta}\!+\!\theta^{\star}).$
@@
In light of Assumption 5, we note that $\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\cdot,\tilde{\theta}_{\text{\tiny AV}})$ is the gradient of a $\mu_{\ell}$-strongly convex function for all $(\tilde{\theta}_{\text{\tiny AV}}+\theta^{\star})\in\Theta$. This observation allows us to establish the following stability properties of the averaged system (27).
@@
$g_{\textsc{r}}\left(\chi,k\right)$ $\coloneq\begin{bmatrix}g_{\textsc{r,1}}\left(\chi_{1},\chi_{2},k\right)\\ g_{\textsc{r,2}}\left(\chi_{1},\chi_{2}\right)\end{bmatrix}\coloneq\begin{bmatrix}\nabla_{3}\tilde{\ell}\left(\chi_{2},\mathbf{1}\sigma(\chi_{2}+x^{\star}),\chi_{1},k\right)-\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\\ \hat{G}^{d}(\chi_{2},\chi_{1})\end{bmatrix}\!$.
@@
The next lemma ensures that the origin is an exponentially stable equilibrium point of (28). To claim this property, we also introduce the set $\mathcal{X}\coloneq\{\chi\in\mathbb{R}^{m+n}\mid\chi_{1}+\theta^{\star}\in\Theta\}$.
@@ IV-E Proof of Theorem 1
The proof consists in finding a uniform ultimate bound to the trajectories of system (17) by using the Lyapunov function $V$ studied in Lemma 4. To this end, by using the notation in (31), we compactly rewrite the original system (17) as
$\xi^{k+1}=F_{\gamma}(\xi^{k},k)-\gamma\mathrm{p}(\xi^{k}),$ (33)
where $\mathrm{p}:\mathbb{R}^{n_{\xi}}\to\mathbb{R}^{n_{\xi}}$ reads as
@@
$\mathrm{p}(\xi)\coloneq\begin{bmatrix}\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\\ \mathrm{p}_{1}(\xi_{2}+x^{\star})+\nabla\phi\left(\xi_{2}+x^{\star}\right)\frac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}(\xi_{2}+x^{\star})\\ 0\end{bmatrix}$.
Let us find a bound for $\left\lVert\mathrm{p}(\xi)\right\rVert$. By definition, we have
@@
$\left\lVert\mathrm{p}(\xi)\right\rVert^{2}$ $=\left\lVert\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)+\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}+\left\lVert\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\right\rVert^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\!\left\lVert\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}+\left\lVert\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}+2\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)^{\top}\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)$
$\hskip 11.0pt+\left\lVert\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\right\rVert^{2}$
@@
$\stackrel{{\scriptstyle(b)}}{{\leq}}\!\left\lVert\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}+\left\lVert\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\right\rVert\left\lVert\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}+2\left\lVert\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)\right\rVert\left\lVert\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\right\rVert\left\lVert\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)\right\rVert$
@@
$\hskip 11.0pt+\left\lVert\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\right\rVert^{2}$
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}\epsilon^{2}N\left(2+3L_{\phi}\right),$
@@
where in $(a)$ we have expanded the first squared norm, in $(b)$ we have used the property that for each $v\in\mathbb{R}$ then $v\leq\lvert v\rvert$ and the Cauchy-Schwarz inequality, while in $(c)$ we leverage the accuracy bounds enforced by Assumption 6, the Lipschitz continuity of each $\phi_{i}(\mathrm{x}_{i})$ (cf. Assumption 2), and the fact that $\left\lVert\mathbf{1}\mathbf{1}^{\top}/N\right\rVert=1$. With this bound at hand, we now consider the Lyapunov function $V$ introduced in Lemma 4 to study the stability properties of the original system (33). Let $a_{5}\coloneq\sqrt{N\left(2+3L_{\phi}\right)}$. Then, by evaluating the increment of $V$ along the trajectories of (33), for all $k\in\mathbb{N}$ and $\xi\in D$, we get
@@
$\stackrel{{\scriptstyle(a)}}{{=}}V\left(F_{\gamma}\left(\xi,k\right),{k+1}\right)-V\left(\xi,k\right)+V\left(F_{\gamma}\left(\xi,k\right)-\gamma\mathrm{p}\left(\xi\right),k+1\right)-V\left(F_{\gamma}\left(\xi,k\right),k+1\right)$
@@
where in $(a)$ we apply the Young’s inequality with parameter $p\in(0,(a_{3}-a_{2}\tilde{a})/(a_{4}a_{5}L_{\ell}))$ considering an arbitrarily chosen $\tilde{a}\in(0,a_{3}/a_{2})$, in $(b)$ rearrange the terms according to the choice of $p$, and in $(c)$ we use the property (32a). Now, we arbitrarily fix $\gamma\in(0,\bar{\gamma})$, define $a_{6}\!\coloneq\!\frac{a_{4}a_{5}}{\tilde{a}}(\frac{L_{\ell}}{p}+\gamma a_{5})$ and bound (35) as
@@
The proof follows by combining (38) with the quadratic bounds (32a) and the trivial inequality $\left\lVert\mathrm{x}^{k}-x^{\star}\right\rVert\leq\left\lVert\xi^{k}\right\rVert$ (see the definition of $\xi$ in (30)), and by setting $\kappa_{1}\coloneq\sqrt{a_{2}/a_{1}}$, $\kappa_{2}\coloneq\sqrt{1-\gamma\tilde{a}}$, and $B\coloneq\sqrt{a_{6}/a_{1}}$.
@@ V Numerical Simulations
In this section, we test DELTA through numerical simulations. We consider an instance of problem (1) in which $N=20$ and
@@
for all $i\in\{1,\ldots,N\}$, where $x_{i}\in\mathbb{R}$, $\sigma(x)\coloneq\sum_{i=1}^{N}\pi_{i}x_{i}/N$. In detail, for all $i\in\{1,\ldots,N\}$, the parameters $\pi_{i},a_{i},b_{i},c_{i}$ are randomly drawn from $[0,1]$ with a uniform probability distribution. Further, for all $i\in\{1,\ldots,N\}$, we pick the entries of $P_{i}\in\mathbb{R}^{2\times 2}$ in the interval $(0,1)$ such that $P_{i}$ is positive definite, while the entries of $v_{i}\in\mathbb{R}^{2}$ and $q_{i}\in\mathbb{R}$ are randomly drawn from $(0,20)$ with a uniform probability distribution. We pick a randomly generated Erdős-Rényi graph with connectivity parameter $p=0.5$. For all $i\in\{1,\ldots,N\}$, we generate $\mathrm{d}^{k}_{x,i}(k)=5\cos(2\pi k/4)$ and $\mathrm{d}^{k}_{\sigma,i}(k)=5\sin(2\pi k/4)$. Each agent $i$ is endowed with a neural network composed by $2$ layers with $300$ neurons equipped with the softplus function. As for the learning problem (4), we consider the cost functions
$\ell_{i}(u_{x,i},u_{\sigma,i},\theta_{i})\!\coloneq\!\frac{1}{2}\!(f_{i}(u_{x,i},u_{\sigma,i})\!-\!\hat{f}_{i}(u_{x,i},u_{\sigma,i},\theta_{i}))^{2}\!\!\!+\!\left\lVert\theta_{i}\right\rVert^{2}\!\!\!,$
for all $i\in\{1,\ldots,N\}$. For what concerns the algorithm parameters, we empirically choose $\gamma=10^{-4}$. Finally, as for the algorithm initialization, for all $i\in\{1,\ldots,N\}$, we pick $x_{i}^{0}$ sampled from a Gaussian distribution, while $\theta_{i}^{0}$ is chosen by considering the well-known Xavier uniform initialization [37]. Fig. 3 shows the evolution over time of the relative cost error $(f_{\sigma}(\mathrm{x}^{k})-f_{\sigma}(x^{\star}))/{|f_{\sigma}(x^{\star})|}$, where $x^{\star}$ is the optimal solution to problem (1). In particular, we compare the relative cost error achieved by Algorithm 1 with the one obtained by an implementation of the Zeroth-Order (ZO) single point method introduced in [22] and with the Distributed Aggregative Gradient Tracking (DAGT) method proposed in [5], where the exact gradients of the local cost functions $f_{i}$ are used. For all simulations we consider the same stepsize and initial conditions.
As predicted by Theorem 1, Fig. 3 shows that the sequence $\{\mathrm{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 converges to a neighborhood of $x^{\star}$. Then, Fig. 4 shows the evolution over time of the estimation error of the descent direction induced both by the neural networks and the tracking mechanism, i.e., $\left\lVert\hat{G}(\mathrm{x}^{k},\mathrm{w}^{k},\theta^{k})-\nabla f_{\sigma}(\mathrm{x}^{k})\right\rVert$, where $\hat{G}(\mathrm{x}^{k},\mathrm{w}^{k},\theta^{k})$ is the estimated descent direction at iteration $k$ and thus it reads as
@@
Figure 3: Comparison between DELTA, and DAGT [5], and DAGT with ZO gradient approximation by [22]
Figure 4: Evolution over time of $\left\lVert\hat{G}(\mathrm{x}^{k},\mathrm{w}^{k},\theta^{k})-\nabla f_{\sigma}(\mathrm{x}^{k})\right\rVert$.
Fig. 5 highlights the timescale separation of Algorithm 1 since it shows that the convergence of the consensus mechanism occurs in a smaller amount of iterations compared to those needed by the optimization and learning parts (cf. Fig. 3 and 4). Finally in Fig. 6 we provide a graphical understanding of the learning capabilities of an agent. In detail, we show the evolution of the learned tangent space to the local cost $f_{i}$ around the current pair $(\mathrm{x}^{k}_{i},\mathrm{w}^{k}_{i}+\phi_{i}(\mathrm{x}^{k}_{i}))$.
Figure 5: Evolution over time of the tracking dynamics. The dashed lines are the global quantities $\sigma(\mathrm{x}^{k})$ (left) and $\sum_{j=1}^{N}\nabla_{2}\hat{f}_{j}(\mathrm{x}^{k}_{j},\sigma(\mathrm{x}^{k}),\theta_{j}^{k})/N$ (right), while the solid lines are the corresponding local agents estimates.

### v3

@@ IV-B Boundary-Layer System Analysis
In this section, we analyze the boundary-layer system associated with (18), obtained by fixing an arbitrary $(\tilde{\theta},\tilde{\mathrm{x}})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ in (18c)-(18d), written in the error coordinates.
@@ IV-C Reduced system analysis
$\hskip 13.0pt+\nabla\phi(\tilde{\mathrm{x}}+x^{\star})\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\hat{G}_{2}(\tilde{\mathrm{x}}+x^{\star},\mathbf{1}\sigma(\tilde{\mathrm{x}}+x^{\star}),\tilde{\theta}+\theta^{\star})$
@@
Using the accuracy characterization (6c) of $\theta^{\star}$ and the optimality of $x^{\star}$ for (1), we point out that the origin is an equilibrium of (25). We then study its stability via the corresponding averaged system, obtained by averaging the time-varying vector field of (25) over an infinite time horizon (see Appendix A-A). To this end, let $\tilde{\ell}_{\mathcal{K}}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ be
$\tilde{\ell}_{\mathcal{K}}(\tilde{x},\tilde{\theta})\coloneq\tfrac{1}{\mathcal{K}}\textstyle\sum_{k=1}^{\mathcal{K}}\ell(\tilde{x}\!+\!x^{\star}\!+\!\mathrm{d}^{k}_{x},\mathbf{1}\sigma(\tilde{x}\!+\!x^{\star})\!+\!\mathrm{d}^{k}_{\sigma},\tilde{\theta}\!+\!\theta^{\star}).$
@@
From Assumption 5 $\nabla_{2}\tilde{\ell}_{\mathcal{K}}(\cdot,\tilde{\theta}_{\text{\tiny AV}})$ is the gradient of a $\mu_{\ell}$-strongly convex function for all $(\tilde{\theta}_{\text{\tiny AV}}+\theta^{\star})\in\Theta$. We thus obtain the following stability properties of the averaged system (27).
@@
$g_{\textsc{r}}\left(\chi,k\right)$ $\coloneq\begin{bmatrix}g_{\textsc{r,1}}\left(\chi_{1},\chi_{2},k\right)\\ g_{\textsc{r,2}}\left(\chi_{1},\chi_{2}\right)\end{bmatrix}$
$\coloneq\begin{bmatrix}\nabla_{3}\tilde{\ell}\left(\chi_{2},\mathbf{1}\sigma(\chi_{2}+x^{\star}),\chi_{1},k\right)-\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\\ \hat{G}^{d}(\chi_{2},\chi_{1})\end{bmatrix}\!$.
@@
Let $\mathcal{X}\coloneq\{\chi\in\mathbb{R}^{m+n}\mid\chi_{1}+\theta^{\star}\in\Theta\}$. The next lemma ensures exponential stability properties of the origin for (28).
@@ IV-E Proof of Theorem 1
The proof consists in finding a uniform ultimate bound to the trajectories of system (17) by using the Lyapunov function $V$ studied in Lemma 4. To this end, by following the notation in (31), we compactly rewrite the original system (17) as
$\xi^{k+1}\coloneq F_{\gamma}^{\text{\tiny or}}(\xi^{k},k)=F_{\gamma}(\xi^{k},k)-\gamma\mathrm{p}(\xi^{k}),$ (33)
with $F_{\gamma}^{\text{\tiny or}}:\mathbb{R}^{n_{\xi}}\to\mathbb{R}^{n_{\xi}}$ and $\mathrm{p}:\mathbb{R}^{n_{\xi}}\to\mathbb{R}^{n_{\xi}}$ which reads as
@@
$\mathrm{p}(\xi)\coloneq\begin{bmatrix}\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\\ \mathrm{p}_{1}(\xi_{2}+x^{\star})+\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}(\xi_{2}+x^{\star})\\ \tfrac{R^{\top}}{\gamma}\left(\phi\left(F_{\gamma,2}^{\text{\tiny or}}(\xi,k)+x^{\star}\right)-\phi\left(F_{\gamma,2}(\xi,k)+x^{\star}\right)\right)\\ \tfrac{R^{\top}}{\gamma}\left(\hat{G}_{2}^{\text{\tiny or}}\left(F_{\gamma}^{\text{\tiny or}}\left(\xi,k\right)\right)-\hat{G}_{2}^{\text{\tiny nom}}\left(F_{\gamma}\left(\xi,k\right)\right)\right)\end{bmatrix}$,
in which, $F_{\gamma,i}(\xi^{k},k)$ and $F_{\gamma,i}^{\text{\tiny or}}(\xi^{k},k)$ denote the $i$-th component of $F_{\gamma}(\xi^{k},k)$ and $F_{\gamma}^{\text{\tiny or}}(\xi^{k},k)$, while $\hat{G}_{2}^{\text{\tiny or}}(F_{\gamma}^{\text{\tiny or}}(\xi,k))$ and $\hat{G}_{2}^{\text{\tiny nom}}(F_{\gamma}(\xi,k))$ represent the function $\hat{G}_{2}$ evaluated respectively at $\left(F_{\gamma,2}^{\text{\tiny or}}(\xi,k)\!+\!x^{\star},\mathbf{1}\sigma\left(F_{\gamma,2}^{\text{\tiny or}}(\xi,k)\!+\!x^{\star}\right),F_{\gamma,1}^{\text{\tiny or}}(\xi,k)\!+\!\theta^{\star}\right)$ and $\left(F_{\gamma,2}(\xi,k)\!+\!x^{\star},\mathbf{1}\sigma\left(F_{\gamma,2}(\xi,k)\!+\!x^{\star}\right),F_{\gamma,1}(\xi,k)\!+\!\theta^{\star}\right)$.
We proceed bounding $\left\lVert F_{\gamma,2}^{\text{\tiny or}}(\xi^{k}\!,k)\!-\!F_{\gamma,2}(\xi^{k}\!,k)\right\rVert^{2}$ as
@@
$\left\lVert F_{\gamma,2}^{\text{\tiny or}}(\xi^{k}\!,k)\!-\!F_{\gamma,2}(\xi^{k}\!,k)\right\rVert^{2}$ $=\big\lVert\mathrm{p}_{1}\!\left(\xi_{2}\!+\!x^{\star}\right)\!+\!\nabla\phi\left(\xi_{2}\!+\!x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\!\left(\xi_{2}\!+\!x^{\star}\right)\big\rVert^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left\lVert\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}+\left\lVert\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)\right\rVert^{2}$
@@
$\hskip 11.0pt+2\mathrm{p}_{1}\left(\xi_{2}+x^{\star}\right)^{\top}\nabla\phi\left(\xi_{2}+x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}+x^{\star}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\epsilon^{2}N\left(1+L_{\phi}^{2}+2L_{\phi}\right)$
$=\epsilon^{2}N\left(1+L_{\phi}\right)^{2},$ (34)
where in $(a)$ we expanded the square norm, while in $(b)$ we used the Cauchy-Schwarz inequality, the Lipschitz continuity of $\phi$ (see Assumption 2), and the accuracy bounds enforced by Assumption 6. Similarly, we compute the following bound
$\left\lVert\mathbf{1}\sigma(F_{\gamma,2}^{\text{\tiny or}}(\xi^{k},k)+x^{\star})-\mathbf{1}\sigma(F_{\gamma,2}(\xi^{k},k)+x^{\star})\right\rVert^{2}$ $=\left\lVert\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\left(\phi(F_{\gamma,2}^{\text{\tiny or}}(\xi^{k},k))-\phi(F_{\gamma,2}(\xi^{k},k))\right)\right\rVert^{2}$
$\leq L_{\phi}^{2}\left\lVert F_{\gamma,2}^{\text{\tiny or}}(\xi^{k},k)-F_{\gamma,2}(\xi^{k},k)\right\rVert^{2}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\epsilon^{2}NL_{\phi}^{2}\left(1+L_{\phi}\right)^{2},$ (35)
where in $(a)$ we used the bound (IV-E). Lastly, we notice that
@@
$\left\lVert F_{\gamma,1}^{\text{\tiny or}}(\xi^{k},k)-F_{\gamma,1}(\xi^{k},k)\right\rVert^{2}=\left\lVert\gamma\mathrm{p}_{\ell}(\xi_{2}+x^{\star},k)\right\rVert^{2}\leq\gamma^{2}\epsilon^{2}N.$ (36)
At this point, $\left\lVert\mathrm{p}(\xi)\right\rVert$ can be bounded as
$\left\lVert\mathrm{p}(\xi)\right\rVert^{2}$ $=\left\lVert\mathrm{p}_{\ell}(\xi_{2}\!+\!x^{\star},k)\right\rVert^{2}$
$\hskip 13.0pt\!+\!\left\lVert\mathrm{p}_{1}\left(\xi_{2}\!+\!x^{\star}\right)\!+\!\nabla\phi\!\left(\xi_{2}\!+\!x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}\!+\!x^{\star}\right)\right\rVert^{2}$
$\hskip 13.0pt\!+\!\left\lVert\tfrac{R^{\top}}{\gamma}\!\left(\phi\!\left(F_{\gamma,2}^{\text{\tiny or}}(\xi\!,k)\!+\!x^{\star}\right)\!-\!\phi\!\left(F_{\gamma,2}(\xi\!,k)\!+\!x^{\star}\right)\right)\right\rVert^{2}$
$\hskip 13.0pt\!+\!\left\lVert\tfrac{R^{\top}}{\gamma}\hat{G}_{2}^{\text{\tiny or}}\left(F_{\gamma}^{\text{\tiny or}}\left(\xi,k\right)\right)-\hat{G}_{2}^{\text{\tiny nom}}\left(F_{\gamma}\left(\xi,k\right)\right)\right\rVert^{2}$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left\lVert\mathrm{p}_{\ell}(\xi_{2}\!+\!x^{\star},k)\right\rVert^{2}$
$\hskip 13.0pt\!+\!\left\lVert\mathrm{p}_{1}\left(\xi_{2}\!+\!x^{\star}\right)\!+\!\nabla\phi\!\left(\xi_{2}\!+\!x^{\star}\right)\tfrac{\mathbf{1}\mathbf{1}^{\top}}{N}\mathrm{p}_{2}\left(\xi_{2}\!+\!x^{\star}\right)\right\rVert^{2}$
$\hskip 13.0pt\!+\!\tfrac{L_{\phi}^{2}}{\gamma^{2}}\left\lVert R\right\rVert^{2}\left\lVert F_{\gamma,2}^{\text{\tiny or}}(\xi^{k}\!,k)\!-\!F_{\gamma,2}(\xi^{k}\!,k)\right\rVert^{2}$
$\hskip 13.0pt\!+\!\tfrac{L_{\sigma}^{2}}{\gamma^{2}}\left\lVert R\right\rVert^{2}\left(\left\lVert F_{\gamma,2}^{\text{\tiny or}}(\xi^{k}\!,k)\!-\!F_{\gamma,2}(\xi^{k}\!,k)\right\rVert^{2}\right$.
$\hskip 13.0pt\!+\!\left\lVert\mathbf{1}\sigma(F_{\gamma,2}^{\text{\tiny or}}(\xi^{k}\!,k)\!+\!x^{\star})\!-\!\mathbf{1}\sigma(F_{\gamma,2}(\xi^{k}\!,k)\!+\!x^{\star})\right\rVert^{2}$
$\left.\hskip 13.0pt\!+\!\left\lVert F_{\gamma,1}^{\text{\tiny or}}(\xi^{k}\!,k)\!-\!F_{\gamma,1}(\xi^{k}\!,k)\right\rVert^{2}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\epsilon^{2}N\!\Big[\!1\!+\!\left(1\!+\!L_{\phi}\right)^{2}\!+\!\tfrac{L_{\phi}^{2}}{\gamma^{2}}\left\lVert R\right\rVert^{2}\left(1\!+\!L_{\phi}\right)^{2}$
$\hskip 13.0pt\!+\!\tfrac{L_{\sigma}^{2}}{\gamma^{2}}\!\left\lVert R\right\rVert^{2}\!\left(\!\left(1\!+\!L_{\phi}\right)^{2}\!\!+\!L_{\phi}^{2}\left(1\!+\!L_{\phi}\right)^{2}\!\!+\!\gamma^{2}\right)\!\!\Big]$
$\eqcolon\epsilon^{2}a_{5}^{2},$
@@
where in $(a)$ we used the Cauchy-Schwarz inequality and the Lipschitz continuity of $\phi_{i}(\mathrm{x}_{i})$ and $\nabla_{2}\hat{f}_{i}$ (cf. Assumptions 2 and 4), while in $(b)$ we used equations (IV-E)-(36). With this bound at hand, we now consider the Lyapunov function $V$ introduced in Lemma 4 to study the stability properties of the original system (33). Then, by evaluating the increment of $V$ along the trajectories of (33), for all $k\in\mathbb{N}$ and $\xi\in D$, we get
@@
$\stackrel{{\scriptstyle(a)}}{{=}}V\left(F_{\gamma}\left(\xi,k\right),{k+1}\right)-V\left(\xi,k\right)$
$\hskip 13.0pt+V\left(F_{\gamma}\left(\xi,k\right)-\gamma\mathrm{p}\left(\xi\right),k+1\right)-V\left(F_{\gamma}\left(\xi,k\right),k+1\right)$
@@
where in $(a)$ we apply the Young’s inequality with parameter $p\in(0,(a_{3}-a_{2}\tilde{a})/(a_{4}a_{5}L_{\ell}))$ considering an arbitrarily chosen $\tilde{a}\in(0,a_{3}/a_{2})$, in $(b)$ we rearrange the terms according to the choice of $p$, and in $(c)$ we use (32a). Now, we arbitrarily fix $\gamma\in(0,\bar{\gamma})$, define $a_{6}\!\coloneq\!\frac{a_{4}a_{5}}{\tilde{a}}(\frac{L_{\ell}}{p}+\gamma a_{5})$ and bound (38) as
@@
[…]
@@ V Numerical Simulations
In this section, we test DELTA through numerical simulations. We consider an instance of problem (1) with $N=20$ agents. In detail, for all $i\in\{1,\ldots,N\}$, we consider the following setup. We define the local cost functions $f_{i}$ as
@@
where $x_{i}\in\mathbb{R}$, $\sigma(x)\coloneq\sum_{i=1}^{N}\pi_{i}x_{i}/N$. Further, the parameters $\pi_{i},a_{i},b_{i},c_{i}$ and the entries of $P_{i}\in\mathbb{R}^{2\times 2}$ are drawn uniformly from $(0,1)$ enforcing $P_{i}$ to be positive definite. The entries of $v_{i}\in\mathbb{R}^{2}$ and $q_{i}\in\mathbb{R}$ are drawn uniformly from $(0,20)$. Then, we set $\mathrm{d}^{k}_{x,i}(k)=5\cos(2\pi k/4)$ and $\mathrm{d}^{k}_{\sigma,i}(k)=5\sin(2\pi k/4)$. We pick local neural networks composed of $2$ layers with $300$ neurons equipped with the softplus function. As for the learning problem (4), we consider the local cost functions
$\ell_{i}(u_{x,i},u_{\sigma,i},\theta_{i})\!\coloneq\!\frac{1}{2}\!(f_{i}(u_{x,i},u_{\sigma,i})\!-\!\hat{f}_{i}(u_{x,i},u_{\sigma,i},\theta_{i}))^{2}\!\!\!+\!\left\lVert\theta_{i}\right\rVert^{2}\!\!\!.$
Further, we pick $x_{i}^{0}$ sampled from a Gaussian distribution, while $\theta_{i}^{0}$ is chosen by considering the Xavier uniform initialization [18]. We empirically tune $\gamma=10^{-4}$ and consider a random Erdős–Rényi graph with connectivity $p=0.5$. Fig. 3a shows the evolution of the relative cost error $(f_{\sigma}(\mathrm{x}^{k})-f_{\sigma}(x^{\star}))/|f_{\sigma}(x^{\star})|$, where $x^{\star}$ is the optimizer of (1). We compare Algorithm 1 with an implementation of the Zeroth-Order (ZO) single-point method of [15] and with the Distributed Aggregative Gradient Tracking (DAGT) method of [20], which uses the exact gradients of $f_{i}$. All simulations use the same stepsize and initial conditions. All neural networks are implemented in TensorFlow, which provides built-in automatic differentiation [2].
As predicted by Theorem 1, Fig. 3a shows that the sequence $\{\mathrm{x}^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 converges to a neighborhood of $x^{\star}$. Then, Fig. 3b shows the evolution over time of the estimation error of the descent direction induced both by the neural networks and the tracking mechanism, i.e., $\hat{G}(\mathrm{x}^{k},\mathrm{w}^{k},\theta^{k})-\nabla f_{\sigma}(\mathrm{x}^{k})||$, where $\hat{G}(\mathrm{x}^{k},\mathrm{w}^{k},\theta^{k})$ is the estimated descent direction at iteration $k$ and thus it reads as
@@
Figure 3: a) Comparison between DELTA, and DAGT [20], and DAGT with ZO gradient approximation by [15]. b) Evolution over time of $\hat{G}(\mathrm{x}^{k},\mathrm{w}^{k},\theta^{k})-\nabla f_{\sigma}(\mathrm{x}^{k})||$
Fig. 4 illustrates the timescale separation in Algorithm 1, showing that the consensus mechanism converges in fewer iterations than the optimization and learning components (cf. Figs. 3a and 3b).
Figure 4: Evolution of the tracking dynamics. The dashed lines are the global $\sigma(\mathrm{x}^{k})$ (left) and $\sum_{j=1}^{N}\nabla_{2}\hat{f}_{j}(\mathrm{x}^{k}_{j},\sigma(\mathrm{x}^{k}),\theta_{j}^{k})/N$ (right). The solid lines are their local agents estimates.
In Fig. 5 we provide a graphical understanding of the learning capabilities of an agent. In detail, we show the evolution of the learned tangent space to the local cost $f_{i}$ around the current pair $(\mathrm{x}^{k}_{i},\mathrm{w}^{k}_{i}+\phi_{i}(\mathrm{x}^{k}_{i}))$.
@@
Finally, we tested the robustness of DELTA to changes in the local $f_{i}$. In particular, at iteration $k=10^{5}$ the parameters of each $f_{i}$ are perturbed by subtracting random values drawn from a uniform distribution in $(0,0.1)$. Figure 6 shows that after this sudden change, the algorithm re-learns the new cost function gradients and converges to a neighborhood of the new optimal solution.
Figure 6: Robustness of DELTA to changes in the local costs. (a) Evolution of the relative cost error. (b) Evolution of the norm of the descent estimation error.